Shift report 20030523 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri May 23 2003 - 11:39:48 EDT

  • Next message: Brendan Fox: "Re: end of brahms+phobos pp run"
    ------------------------------------------------------------
    Date: Fri May 23 11:39:48 2003, Shift: 08:00-16:00
    Supervisor: db, with: db + KH
    Log book#14 pages 115 - 117
    
    Shift Summary:
    ------------------------------------------------------------
    Store started at the beginnning of the shift. Tookrun 9026 for 10k trig 2 and 27k trig 3
    Run 9027 was stopped early since we were told this was the last store.
    run 9028 47k 1/8B on FS
    run 9029 38k 1/4B on FS  Store died.
    Put chambers in park. Told E Baker to secure flammable gas today
    Magnets placed in standby.
    I have not truned off pmt since people may want to do things
    TURN OUT THE LIGHTS THE MUSIC IS OVER.
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Fri May 23 2003 - 11:40:23 EDT